OpenStudy (anonymous):

list the missing value after completing the square for this equation. x^2+10x-7=0 x=-5+4 sqrt __?

OpenStudy (anonymous):

it says 128 is wrong???

OpenStudy (anonymous):

no it is \[5\pm4\sqrt{2}\]

OpenStudy (amistre64):

there is a way to dbl chk answers instead of blindly filling in any one you get ;)

OpenStudy (anonymous):

lol yeah im just trying to finish my last worksheet so i can get my credit for this class

OpenStudy (amistre64):

understandable ☺☻☺

OpenStudy (anonymous):

like so \[x^2+10x=7\] \[(x+5)^2=7+25=32\] \[x+5=\pm\sqrt{32}\] \[x=-5 \pm\sqrt{32}\] \[x=-5\pm 4\sqrt{2}\]

OpenStudy (anonymous):

last step because \[\sqrt{32}=\sqrt{16\times 2}=\sqrt{16}\times \sqrt{2}=4\sqrt{2}\]